### v1

# Student-AI Interaction: A Case Study of CS1 students

Abstract.
The new capabilities of generative artificial intelligence tools (Generative AI), such as ChatGPT, allow users to interact with the system in intuitive ways, such as simple conversations, and receive (mostly) good-quality answers. These systems can support students’ learning objectives by providing accessible explanations and examples even with vague queries. At the same time, they can encourage undesired help-seeking behaviors by providing solutions to the students’ homework. Therefore, it is important to better understand how students approach such tools and the potential issues such approaches might present for the learners.
In this paper, we present a case study for understanding student-AI collaboration to solve programming tasks in the CS1 introductory programming course. To this end, we recruited a gender-balanced majority non-white set of 15 CS1 students at a large public university in the US. We observed them solving programming tasks. We used a mixed-method approach to study their interactions as they tackled Python programming tasks, focusing on when and why they used ChatGPT for problem-solving. We analyze and classify the questions submitted by the 15 participants to ChatGPT. Additionally, we analyzed user interaction patterns, their reactions to ChatGPT’s responses, and the potential impacts of Generative AI on their perception of self-efficacy.
Our results suggest that in about a third of the cases, the student attempted to complete the task by submitting the full description of the tasks to ChatGPT without making any effort on their own. We also observed that few students verified their solutions. We discuss the results and their potential implications.
1. Introduction
Many interactive learning environments (ILEs) offer on-demand help, which refers to information actively solicited by the learner and provided by the system, intended to positively influence learning (Aleven et al., 2003). While effective help-seeking behavior in ILEs is linked to improved learning outcomes, many learners do not effectively utilize available help resources(Aleven and Koedinger, 2000; Luckin et al., 1999). This raises concerns that ILEs may not reach their full potential unless we find ways to help students use these support tools better. Additionally, since seeking help is a crucial skill that affects learning in many situations(Nelson-Le Gall, 1981; Newman, 2023; Aleven and Koedinger, 2001), designing ILEs to encourage effective help-seeking behaviors could greatly enhance their educational value(Aleven and Koedinger, 2001; Gräsel et al., 2000).
Generative AI can be considered an interactive learning environment (ILE), as they can provide on-demand help to students, such as providing explanations of concepts and offering examples. They play a crucial role as interactive components within broader educational settings. They contribute to interactive learning by providing immediate responses and facilitating learning interactions in real time. With the advance of Generative AI tools such as ChatGPT and Copilot, they have become integral tools for many students, particularly those in introductory programming courses(Prather et al., 2023a). They can assist students by providing code suggestions, explaining programming concepts, identifying and resolving errors, and generating code documentation (Kazemitabaar et al., 2023b; Denny et al., 2024c). Their widespread adoption signifies a major shift in how programming is learned and practiced. However, this increasing reliance on Generative AI tools presents significant challenges. One primary concern is that students might become overly dependent on these tools, potentially hindering their ability to develop fundamental programming skills (Prather et al., 2024). Additionally, AI-generated solutions might lead to academic dishonesty or reduced problem-solving capabilities among students (Becker et al., 2023). This dichotomy has sparked a debate among educators: while some argue against the use of Generative AI in learning due to these risks, others advocate for its judicious use to enhance learning outcomes (Lau and Guo, 2023).
Despite various studies exploring the impact of Generative AI tools like GitHub Copilot on productivity (Ziegler et al., 2022a; Kazemitabaar et al., 2023a) and learning outcomes (Tan et al., 2024; Margulieux et al., 2024), there remains a critical gap in understanding how novice programmers interact with these tools without any limitations on using Generative AI, where they can freely utilize AI-driven assistance and guidance. Understanding students’ level of trust in these tools, which impacts their adoption and learning outcomes (Amoozadeh et al., 2024, 2023), and analyzing their interactions while using Generative AI for programming tasks is crucial. These insights can help us identify behaviors, actions, decisions, and responses while utilizing Generative AI tools. By modeling their behaviors and identifying challenges, we can develop systems that provide effective AI-assisted learning, thereby enabling students to learn programming concepts more efficiently and achieve better educational outcomes.
To address this gap, our study investigates the interaction of novice programmers with ChatGPT while solving programming tasks. We conducted a mixed-method study involving 15 CS1 students using a specially designed VSCode plugin that integrates ChatGPT directly into their coding environment. This setup allows us to observe their natural use of Generative AI assistance without external constraints. Our study aims to understand how, when, and why students use ChatGPT, and what strategies they employ to integrate Generative AI into their problem-solving processes.
This paper aims to better understand how students use ChatGPT during programming tasks, revealing diverse usage patterns and their effectiveness. In particular, we seek to answer the following research questions:
  • RQ1:
    How frequently and at what stages do students use ChatGPT while solving programming tasks?
  • RQ2:
    How do students interact with ChatGPT while solving a programming task?
  • RQ3:
    What strategies do novices employ when interacting with ChatGPT?
Our findings highlight that participants, such as P2, extensively interacted with Generative AI yet faced challenges in achieving accurate solutions. Others adopted approaches including step-by-step prompting or hybrid approaches that combine independent coding and Generative AI support. The acceptance of the Generative AI responses varied, from full adoption of the Generative AI’s response to using Generative AI to find answers to the queries, or simply comparing the Generative AI responses to their own solutions. Collaboration success rates showed that Generative AI facilitated correct answers in 65% of the cases, while 34% remained unresolved despite using Generative AI.
2. Related Work
2.1. Exploring the Impact of GenAI in Programming
The recent widespread deployment of generative AI programming assistants has motivated much research on the use of these tools. Researchers have conducted empirical studies (Leinonen et al., 2023; Hellas et al., 2023; Sarsa et al., 2022) to evaluate the quality of code and explanations that GenAI programming assistants can generate. Others have explored the usefulness of LLM-based programming tools through user studies (Ziegler et al., 2022b; Vaithilingam et al., 2022; Barke et al., 2023; kazemitabaar2023; Xu et al., 2022; Mozannar et al., 2022; Ross et al., 2023; Srinivasa Ragavan and Alipour, 2024). Notably, Vaithilingam (Vaithilingam et al., 2022) compared the user experience of GitHub Copilot with traditional auto-complete and found that programmers faced more frequent difficulties in completing tasks with Copilot, although there was no significant impact on task completion time. Barke (Barke et al., 2023) took a step further to understand how programmers interact with code-generating models using a grounded theory analysis, identifying two primary modes of interaction: acceleration, where GenAI is used to speed up code authoring in small logical units, and exploration, where GenAI serves as a planning assistant by suggesting structure or API calls. While these studies have contributed to our understanding of how developers use GenAI developer tools, they mainly focused on evaluating the effectiveness of such tools and did not directly address developers’ perceptions of them. Some studies, such as those by Liang (Liang et al., 2023) and Ziegler (Ziegler et al., 2022b), have also focused on how developers perceive these tools. For instance, Liang (Liang et al., 2023) conducted a survey involving 410 developers to investigate the usability challenges associated with GenAI programming assistants, finding that although developers appreciate their auto-complete capabilities, they often face challenges in controlling the quality of code generation and have concerns about potentially infringing on intellectual property. Although these studies provide valuable insights on how users perceive the GenAI tools, the insights are primarily applicable to professional programmers.
2.2. The Role of AI in Education
Recent studies have started to explore these GenAI tools’ applications in educational contexts. For example, Denny et al. (Denny et al., 2024b) have identified desirable characteristics for GenAI teaching assistants in programming education. They emphasize the importance of features that provide instant, engaging support while allowing students to maintain autonomy in their learning process. Liu et al. (Liu et al., 2024) investigated the integration of GenAI tools in the CS50 course at Harvard University, demonstrating how generative GenAI can significantly enhance the learning experience for introductory programming students. Their approach involved developing and deploying a suite of GenAI-based tools designed to emulate a 1:1 teacher-to-student ratio, thereby providing personalized, real-time support to students. These tools, including the CS50 Duck, a virtual rubber-duck debugging assistant, were well-received by students, who reported that the GenAI tools made them feel as though they had a personal tutor available at all times. Prather et al. (Prather et al., 2023b) have recently published a series of studies examining the usability and interaction challenges faced by novice programmers using GenAI tools like GitHub Copilot. Their research highlights the dual nature of such tools: while they can significantly accelerate the coding process and aid in overcoming programming blocks, they also introduce unique cognitive and metacognitive difficulties. Students often found it ”weird” how accurately Copilot predicted their needs, which led to a mix of fascination and dependency concerns. The study also identified new interaction patterns such as ”shepherding” and ”drifting,” reflecting the nuanced ways novices attempt to guide the GenAI or get misled by it. These insights underscore the importance of designing GenAI tools that support learning while promoting independent problem-solving skills and managing over-reliance.
2.3. Interaction Patterns of Novice Programmers Using LLMs
Kazemitabaar et al. (Kazemitabaar et al., 2023b) conducted a thematic analysis of novice learners using an LLM-based code generator in a self-paced Python course. They identified distinct coding approaches among the learners, noting that the Hybrid approach, which combined manual coding with LLM assistance, produced the best outcomes. However, the study also highlighted signs of over-reliance on LLMs, such as copying LLM output without changes, and positive self-regulation behaviors, like adding code to verify LLM output. Nguyen et al. (Nguyen et al., 2024) conducted a large-scale, multi-institution study that explored the challenges faced by near-novices when interacting with Code LLMs. Their study identified barriers such as difficulties in expressing problem understanding and using appropriate coding terminology. Prather et al. (Prather et al., 2024) further investigated the impact of generative AI tools on novice programmers’ metacognitive awareness and problem-solving strategies, highlighting the potential widening gap between well-prepared and under-prepared students in the era of generative AI.
2.4. Assessing Student Self-Efficacy in AI-Enhanced Learning
Self-efficacy, or the belief in one’s capabilities to achieve a goal or an outcome, is a crucial factor in students’ learning processes. Recent studies have examined how interactions with GenAI tools influence students’ self-efficacy in programming. Tankelevitch et al. (Tankelevitch et al., 2024) discussed the metacognitive demands imposed by GenAI systems, highlighting the need for tools that support students in monitoring and controlling their learning processes. Xue et al. (Xue et al., 2024) investigated the impact of ChatGPT on introductory programming students, finding that the AI tool can enhance students’ self-efficacy by providing immediate feedback and support, thereby reducing the anxiety associated with complex problem-solving tasks. Furthermore, Denny et al. (Denny et al., 2024b) emphasized the importance of designing AI teaching assistants that not only assist with immediate problem-solving but also help students develop long-term self-efficacy by encouraging independent learning and critical thinking.
3. Methodology
The methodology of this study is geared towards examining how CS1 students interact with generative AI tools in a safe environment without limitations on using Generative AI. We specifically chose to implement a ChatGPT plug-in within the VSCode environment. Using ChatGPT within VSCode aims to reduce the cognitive load associated with switching between different environments, thereby allowing participants to remain within a single, cohesive coding environment.
Tankelevitch et al. (Tankelevitch et al., 2024) research highlights the metacognitive demands placed on users when interacting with generative Generative AI systems. These demands include the need for self-awareness of task goals, task decomposition, and evaluating Generative AI-generated outputs (Tankelevitch et al., 2024). This design choice was influenced by studies that suggest minimizing context switching can enhance user focus and reduce metacognitive demands (Tankelevitch et al., 2024). Within this setup, users encounter a question box where they can freely input prompts without any restrictions on the number or type of questions, as shown in Figure 1 ”Ask a question”.
We recruited 15 CS1 students to participate in a user study supervised by one of our authors in an office setting. Each participant was offered bonus points for their involvement. We considered offering bonus points because we did not want to put any pressure on them. We also clarified that their performance would not be judged by others to ensure that the participation was non-coercive. Participants were given the option to utilize Generative AI assistance at their discretion or complete tasks independently.
3.1. Survey Questions
We asked participants to complete a survey before beginning the programming tasks, referred to as the pre-study, and another survey, known as the post-study, after completing the programming tasks. The pre-study and post-study included questions on self-efficacy (Pintrich et al., 1991). Our purpose is to compare participants’ levels of self-efficacy before and after programming, particularly among those who used Generative AI to solve programming tasks. We calculated the mean self-efficacy score for each participant to compare their levels before and after the study. The pre-study was divided into four sections, whereas the post-study solely comprised self-efficacy and demographic questions.
3.2. Participants
In this study, we enrolled a total of 15 participants, comprising 8 females and 7 males. Among them, two participants were first-generation students. Within this group of 15 students, 13 were freshmen, one was a junior, and one was a post-baccalaureate student. All participants were English speakers as their first or second language, and we did not encounter any reported difficulties in explaining concepts or instructions in English. Table 1 shows the participants’ demographics.
Participant Age Gender Education-level Racial/Ethnic First-generation
P1 19 Female Freshman African American No
P2 18 Female Freshman African American No
P3 19 Female Freshman African American No
P4 19 Male Freshman Asian No
P5 18 Female Freshman Asian No
P6 19 Female Freshman Asian Yes
P7 19 Male Freshman African American No
P8 18 Male Freshman Latino Yes
P9 22 Female Postbacc African American No
P10 18 Male Freshman Asian No
P11 19 Female Freshman White No
P12 19 Male Freshman African American No
P13 18 Male Freshman African American No
P14 20 Male Junior Latino No
P15 19 Female Freshman African American No
Table 1. Demographic Information of Participants
3.3. Protocol
To ensure ethical compliance, the informed consent form for this research study was created in adherence to the guidelines set by the Institutional Review Board (IRB) at a large public university in the US. The informed consent form provided participants with essential information about the study, such as its voluntary nature, the estimated time required (here, no more than 60 minutes), and the assurance of confidentiality. It emphasized that no identifiable information would be collected or stored, and participants were given the option to provide their email addresses separately for future study purposes. The form highlighted that there were no expected risks associated with participating in the research and encouraged participants to ask any questions they had before deciding to participate.
3.4. Study Procedure
We provided participants with three programming tasks in an office within the computer science department during the spring semester of 2024. The programming tasks were tailored to the complexity appropriate for a CS1 level and covered topics taught by the instructor throughout the semester preceding the study. Each participant had one hour to respond to the three programming tasks, utilizing the Python programming language that they had learned during the semester. Prior to commencing programming, we offered instructions on how to utilize the plug-in for inputting prompts. It was emphasized that no other Generative AI or aids were permitted during this time. These tasks were designed to encompass several fundamental programming concepts. The first task involved identifying and extracting even numbers from a given set, requiring an understanding of conditional statements and loops for iteration. The second task introduced input/output operations and string manipulation, challenging users to read input from the user and reverse the order of characters. Lastly, the third task delved into the concept of functions and sequences, prompting users to create a function that generates the first n values of the triangular numbers sequence. Each task touched upon recursion or iterative approaches, emphasizing algorithmic thinking and problem-solving skills. Collectively, these tasks addressed core programming concepts such as conditionals, loops, functions, input/output operations, and algorithm design, providing participants with a comprehensive understanding of essential programming principles.
Figure 1. Interface of the VS Code plug-in used in the study: ChatGPT Response and Participant Code
3.5. Data Instrumentation
The plug-in (Nam et al., 2024) we used for this study logs detailed user interactions within the text editor, including all changes to text selections and document modifications, along with the precise timestamps of these events. It also records the prompts users write and the responses generated by ChatGPT. The plug-in interface provides a prompt-writing box on the left side of VSCode, allowing users to access ChatGPT by writing prompts. Additionally, users can see Generative AI responses and their code in the same window, making it easy to compare the Generative AI responses with their code. This setup ensures a focused environment where users are less tempted to use other AIs. Figure 1 shows the interface of the plug-in.
3.6. Data Analysis
3.6.1. Qualitative.
To address our research inquiries, we conducted both qualitative and quantitative analyses of the data. Two authors independently coded the data for three programming tasks, each varying in difficulty levels and topics, utilizing the initial codebook. Subsequently, they engaged in discussions to review the initial coding outcomes, resolve any conflicts, and enhance the codebook further. With the identification of each new code, we reviewed all prior comments for possible adjustments.
The codebook provides a clear breakdown of how participants interacted with Generative AI prompts throughout various stages of programming tasks. It categorizes prompts based on their types and how participants engaged with them, whether standalone or as follow-up interactions. Additionally, it outlines how participants responded to errors during programming and their preferences for resolving them, including using Generative AI responses or ignoring errors altogether. Moreover, the book details the extent to which participants accepted AI-generated solutions, from copying entire responses to incorporating only certain parts or using them for guidance. In essence, the codebook served as a guide for understanding participant-AI interactions during programming activities in educational settings. The codebook includes such thematic coding of responses allowing us to draw meaningful insights about students’ perspectives and shed light on the complex range of attitudes and emotions surrounding Generative AI systems in the context of programming.
3.6.2. Quantitative.
We conducted two surveys, one administered before the programming session and another after, to assess participants’ level of self-efficacy. Participants rated their self-efficacy levels using Likert scales, indicating their agreement with statements related to their programming abilities. These Likert scores provided quantitative data to measure changes in participants’ self-efficacy before and after engaging in the programming tasks.
4. Results
Fifteen CS1 students participated in our students, and we asked each of them to do their best to complete three programming tasks. Participants P3, P5, P6, P9, P12, and P14 solved the three programming tasks successfully. All participants except P3 and P15 had at least one correct answer for the three tasks. Notably, participants P1, P4, P12, and P13 did not complete their last tasks, and P13 also did not complete their second task. Overall, we analyzed 40 solutions submitted by participants, from which 29 solutions were created with assistance from the Generative AI plug-in provided, and 11 without Generative AI.
4.1. Frequency and Adoption Time of Generative AI
We assigned one point to each user request to track how many times users wrote prompts using Generative AI. Table 2 displays the frequency of participants using the Generative AI through ChatGPT plug-in across the three programming tasks in our study, illustrating the different levels of dependence on Generative AI assistance among participants.
Participant Tasks Completed Correct Using Generative AI #Times Generative AI
P1 Condition, String, Function Y, Y, N W, W, N N, Y, N 0, 1, 0
P2 Condition, String, Function Y, Y, Y W, W, C Y, Y, Y 9, 3, 1
P3 Condition, String, Function Y, Y, Y C, C, C N, N, Y 0, 0, 2
P4 Condition, String, Function Y, Y, N C, W, N N, N, N 0, 0, 0
P5 Condition, String, Function Y, Y, Y C, C, C N, Y, Y 0, 1, 10
P6 String, List, Array Y, Y, Y C, C, C N, Y, N 0, 5, 0
P7 String, List, Array Y, Y, Y C, C, W Y, Y, Y 2, 1, 4
P8 String, List, Array Y, Y, Y C, C, W N, N, Y 0, 0, 3
P9 String, List, Array Y, Y, Y C, C, C Y, Y, Y 1, 1, 1
P10 String, List, Array Y, Y, Y C, C, W Y, Y, Y 1, 2, 1
P11 String, List, Array Y, Y, Y W, C, C N, Y, Y 0, 1, 1
P12 String, List, Array Y, Y, N C, C, N Y, Y, N 3, 1, 0
P13 String, List, Array Y, N, N W, N, N Y, N, N 3, 0, 0
P14 String, List, Array Y, Y, Y C, C, C Y, Y, Y 1, 1, 1
P15 String, List, Array Y, Y, Y W, W, W Y, Y, Y 1, 1, 1
Table 2. Usage and Problem Solving by Participants: This table details participants’ problem-solving activities. “Participant” identifies participants, “Tasks” shows the main concepts in the tasks attempted, and “Completed” indicates if the task has been completed by the participant (‘Y’ for yes, ‘N’ for no). “Correct” shows if the submission was correct (‘C’), wrong (‘W’), or not attempted (‘N’). ”Using Generative AI ” indicates Generative AI use (‘Y’ for yes, ‘N’ for no), and “Times Generative AI” shows the number of times Generative AI used.
Table 2 shows P2 made more extensive use of Generative AI compared to other participants. P2 used ChatGPT nine times to solve Q1 but was unable to arrive at a correct solution. This instance shows that despite multiple interactions with Generative AI, the user was not able to solve the task. Participants’ strategies in using Generative AI and their inherent problem-solving abilities both play crucial roles in their success. Some participants, like P3 and P6, achieved correct answers (’C’) without relying on Generative AI. P3, for instance, solved all their tasks correctly without using Generative AI for two out of three programming tasks, indicating that some participants may not need Generative AI to perform well.
As participants did not have any limitations on using Generative AI in terms of when and how many times it was used, we observed that users typically engaged with generative AI at three different stages.
At the beginning of the programming. During the early stages of the programming, participants in 18 tasks turned to ChatGPT right away, without making any initial attempts on their own. Typically, they would type in the full description of the task or paraphrase it. P9, P10, P14, and P15 started using Generative AI right from the beginning, asking ChatGPT to help complete the entire task for them for all three programming tasks.
Middle of programming. In this study, for n=9 questions, Generative AI users sought assistance of Generative AI in the middle of their programming. Initially coding manually, they resorted to using Generative AI due to encountering errors with invalid code or programming concepts.
After solving the problem. Interestingly, two participants (P3 and P5) employed Generative AI even after they had already arrived at the correct answer on their own. P3, for example, used Generative AI to validate she understood the problem description correctly by inquiring about the distinction between ’i’ and ’n’ in the triangular numbers sequence. P5 reached the correct solution on their own first, but still sought Generative AI’s solution for the entire question, seemed that they attempted to verify the correctness of their solution by comparing their solution with the Generative AI’s responses. Table 3 shows when Generative AI users utilized Generative AI.
User When (Generative AI Used) Approach
Q1 Q2 Q3 Q1 Q2 Q3
p1 - Middle - - Hybrid/partial -
p2 Early Middle Early Step by step Hybrid/partial Full description
p3 - - After-Solving - - Step by step
p4 - - - - - -
p5 - After-Solving Early - Full description Step by step
p6 - Middle - Themselves Hybrid/partial -
p7 Middle Middle Middle Hybrid/partial Hybrid/partial Hybrid/partial
p8 - - Middle - - Hybrid/partial
p9 Early Early Early Full description Full description Full description
p10 Early Early Early Full description Full description Full description
p11 - Early Early - Full description Full description
p12 Early Middle - Full description Hybrid/partial -
p13 Middle - - Step by step - -
p14 Early Early Early Full description Full description Full description
p15 Early Early Early Full description Full description Full description
Table 3. Summary of Generative AI Usage and Approaches by Participants: The table summarizes the Generative AI usage and approaches taken by users for different programming questions (Q1, Q2, Q3). The first set of columns under ”When (Generative AI Used)” indicates the timing of Generative AI usage classified as ”Early”, ”Middle”, ”After-Solving”, or ”-” if not used. The second set of columns under ”Approach” details the approach taken by participants, such as ”Hybrid/partial”, ”Step by step”, ”Full description”, and ”-” if not used.
4.2. Users Actions
We identified six common actions users performed when using Generative AI, namely, reading, thinking, writing code, modifying code, prompting, and debugging. Table 4 describes these actions.
Table 4. Description of six common actions users performed when using Generative AI user actions.
Action Description
Reading Reading the question before starting to write code
Thinking Thinking about and confirming a suggestion
Writing Code Writing code by themselves or copying Generative AI response
Modifying Code Editing existing code
Prompting Asking questions from Generative AI
Verification and Debugging Running code and finding errors
The actions described in Table 4 varied among participants, with some repeating actions multiple times and others performing them only once. Figure 2 shows the actions of two participants. Participant P14 consistently followed a straight path across all three programming questions, performing each action once. In contrast, P7 in question 4 repeated certain actions such as running, debugging, prompting, modifying, and running again. Overall, we observed 17 instances of straight actions and 23 instances of repeated actions.
(a) Refer to caption
(b) Refer to caption
Figure 2. Actions of participants: (a) Linear Actions of P14 for all questions, (b) Non-linear Actions of P7 for Question 3
4.3. Approach in Interacting Generative AI
We categorize users’ approaches to Generative AI based on prompts. The table 3 shows how users approach Generative AI. We identified three strategies among users: full description, step-by-step, and partial. In the full description, the user copies the full description of the task into the plugin. In step-by-step, participants broke the question down into structured subgoals and used Generative AI to solve the problem step by step. We also observed a hybrid or partial approach, where participants tackled some parts of the question independently while seeking assistance from Generative AI for some other parts.
Full description
In total, we have n=17 approaches to Generative AI that utilized the entire of question in their prompt. P2 employed a full description of question 3, while for questions 1 and 2, their approach was step-by-step and Hybrid/partial. P5 only utilized the full description for question 2, although they already had the correct answer before resorting to Generative AI. P9, P10, P11, P14, and P15 initially utilized Generative AI by transcribing the full description of the question without writing any meaningful lines of code at the beginning of their programming process. Using a full description of the question does not necessarily guarantee that they will be able to find the correct answer. Among the participants who relied on full descriptions, P10 for Q3, P15 for Q1, Q2, and Q3 were unable to find the correct answers. The finding in Table 3 reveals that the full description approach is commonly used in early programming for most cases.
Step-by-step
Some users attempted to approach the questions step-by-step, breaking them down into various sub-goals. P2, P3, P5, and P13 adopted a similar Generative AI approach to some of the problems. For example, user P5 used Generative AI for the first time to inquire, ’How to round numbers in Python?’ and then followed up with a second query, “How to round numbers in Python only from the tenths place?”. Similarly, P13 used Generative AI three times for their first question. The first inquiry was, “In Python, can you turn an integer into a list?” The second query was about arrays: “What is an array?” Lastly, the user asked: “Given an integer n, produce the array: 0123…n, 1234…0, n012…n-1?.”
Hybrid/Partial
Some participants attempted to write programs independently, but relied on Generative AI for debugging to resolve errors or others. We have identified n=8 approaches that we categorize as Hybrid or Partial. For example, User P2 asked Generative AI, “Fix my code to work.” User P6 encountered an error and copied the message: "TypeError: ’int’ object is not subscriptable".
4.4. Prompt Category
We analyzed the prompts that users submitted to ChatGPT each time they engaged with it. A total of 15 participants wrote 60 prompts. We categorized the prompts based on their relation to questions. The categories include Coding Concepts in Python, Program Logic, Debugging, and entire of Question.
Coding Concepts in Python.
In our study, nearly a third of the prompts (19 out of 60) were related to coding concepts. Seven users sought assistance from Generative AI specifically for coding concepts, indicating their understanding of logic but their need for coding support. Notable inquiries include P2’s query, “How to append to a list?”, P5’s question, “How to round numbers in Python?”, P6’s curiosity about “Python built-in function for arranging list in increasing order”, and P7 asking, “In my program, I am trying to move the negative numbers to the front. Could I initialize an empty list first?”
Program Logic.
In 10 prompts, participants utilized Generative AI because they found it challenging to determine the sequence of steps or instructions needed to accomplish a specific task or goal within the program. For instance, P2 asked about ”a function to return the first n values of the triangular number sequence starting from 1.” P12 needed help trying to reverse a list. P7 inquired, ”I am trying to produce the array 0, 1, 2, 3 all the way to n, meaning it could also be n, 0, 1, 2. Is my approach correct so far?”
Debugging.
Some novices turned to Generative AI for error resolution, indicating that their prompts typically revolved around the question: “How can I solve this error?” In total, 11 prompts were utilized Generative AI for error resolution. Often, users would simply copy the error message from the console and ask ChatGPT for help. P6 copied and pasted the code, asking, “What is wrong with this code?” P10 asked, “Why does the code not work?” P6 specifically inquired about the error received: “TypeError: ‘int’ object is not subscriptable.”
Entire of Problem description
Some Generative AI users simply copied the question descriptions, which we considered as their entire prompt in the question category. A total of 20 out of 60 prompts fell into this category. In some cases, such as P11, asking the full questions 2 and 3 resulted in receiving the entire answer with just one prompt.
4.5. Users’ reaction to the Generative AI responses
We identify various user actions upon receiving responses from ChatGPT, which we categorize as acceptance categories. These actions reflect users’ reactions to Generative AI answers, and this collaboration can result in either successes or failures. These categories are based on the prompts and their reactions to the Generative AI’s answers.
4.5.1. Acceptance Category
Complete use
Occasionally, participants completely accept the Generative AI’s response in its entirety. We have $n=18$ prompts where users accept Generative AI answers based on these prompts. For example, P10 formulated a prompt for Question 3 “Given an integer n, produce the array” and accepted the Generative AI’s solution without any modification. Similarly, P7, when addressing the same question, queried “I am trying to produce the array 0,1,2,3 all the way to n, meaning it could also be n,0,1,2 is my correct so far?” fully embraced the Generative AI’s response.
Selective use
In this scenario, some participants tried to understand the Generative AI response and figure out how to implement it in their own way. There are $n=23$ prompts where users get ideas from Generative AI responses to write their own code. For instance, P5 asked: “How to round numbers in Python only from the tenths place” and the Generative AI response was “To round numbers in Python only from the tenths place, you can use the round() function”, followed by an example of how round() function works. The user completed the task on their code.
Retry
Using recorded videos of facial expressions and eye movements, we identified cases in which participants spent an extended amount of time reading and (most likely) contemplating Generative AI’s responses. We observed six participants visibly contemplating the Generative AI’s response to 19 prompts. These participants read the response but did not copy the code or write it in their own way. In such cases, users attempted to write another prompt, which may indicate their rejection of the Generative AI solution. P12’s initial query was “I need help trying to reverse a list.” After reading the Generative AI response, the user attempted to write another prompt, “Help me turn a sequence of numbers into a list.” Once again, the user read the Generative AI’s response. Subsequently, P12 revised his prompt and wrote: “Without using a built-in function, help me turn a sequence of numbers into a list.” After this prompt, he began to write code in their own way.
4.6. Correctness of Generative AI-assisted solution
To assess the frequency of successful collaboration facilitated by Generative AI, we used manual inspection and testing of the sample inputs to evaluate the correctness of the submissions. We considered all the cases that the participants used Generative AI at any stage in problem-solving. Interestingly, among the $n=29$ solutions in which participants used Generative AI, 19 solutions ($65\%$) were correct and the rest (35%) were incorrect. Most of the problems stem from incorrect copy-paste from the ChatGPT plugin, and the rest were from the difficulty in integrating portions of the ChatGPT responses into the participants’ current solution.
4.7. Participants’ Self-efficacy
To evaluate the potential impact of Generative AI on self-efficacy, we conducted pre-study and post-study self-efficacy surveys. Unfortunately, for the first five participants (P1 to P5), we did not capture the pre-study self-efficacy data. Figure 3 presents the participants’ perception of self-efficacy before and after using Generative AI. Figure 3(a) shows the distribution of perception of self-efficacy before and after the study. Given the small number of observations, we cannot draw any statistically-sound conclusion about any difference between the distributions.
(a) Distribution of Pre-study and Post-study self-Efficacy scores
(b) Pre- vs. Post-study self-efficacy scores by participant
Figure 3. Self-efficacy of participants
Figure 3(b) compares the self-efficacy of the participants before and after the experiment. It shows that P6 experienced a noticeable increase in self-efficacy, from a self-efficacy score of 4.00 in pre-study to 4.75 after the study, indicating a potential positive impact of using Generative AI on self-efficacy. P6 completed all three programming tasks successfully and demonstrated a mix-model behavior by independently addressing two tasks and using a hybrid approach with Generative AI for one task. The increase in self-efficacy scores suggests that employing Generative AI was effective in improving her perception of self-efficacy.
In contrast, self-efficacy in P9 decreased from 2.58 before the study to 1.75 after the study. Although P9 completed all programming tasks successfully, she copied the full description to the plug-in, asked ChatGPT for the answer, and pasted the results as the submission. Similarly, this occurred for P13, who began with a pre-study self-efficacy score of 3.92, however, the post-study mean score dropped to 3.00. He submitted an incorrect solution for one task and did not finish the other two tasks, she used Generative AI in a step-by-step approach.
5. Threats to Validity
Replicability Can others replicate our results? Because Generative AI and its usage and acceptance are rapidly emerging, we do not know if the results of this study will be replicable due to changing student behaviors and rapid advances in Generative AI. We encourage other researchers to replicate our study for different populations to check for emerging behaviors of students in using Generative AI.
Internal Did we skew the accuracy of our results with how we collected and analyzed data? We used open coding to analyze the data and two authors reviewed all the data to make sure that the coding of videos adhered to the codebook that was created.
External Do our results generalize? Because our study has a small sample size (n=15) relative to the overall student population, it is not possible to generalize to all students’ behavior. However, we attempted to have diversity based on gender. The bias of drawing conclusions from self-selection bias remains a threat to validity.
6. Discussion and Concluding Remarks
In this section, we discuss the potential implications of our findings for computing education research and practice.
6.1. Interactions with Generative AI
In this study, our objective was to investigate how CS1 students utilize ChatGPT for programming questions to assess whether Generative AI serves as a help-seeking tool for beginners in programming. As observed in other recent studies  (Kazemitabaar et al., 2023b; Prather et al., 2024), our findings indicate that a large number of Generative AI users rely on providing full descriptions of programming questions to find solutions without making sufficient effort on their own, even under supervision. This trend aligns with the patterns of over-reliance on LLMs identified by Kazemitabaar et al.  (Kazemitabaar et al., 2023b), particularly among novices using the ’AI single prompt’ approach, which resulted in lower performance on subsequent tasks. The observed behavior raises concerns about the potential overreliance on Generative AI in educational settings, where students might increasingly rely on Generative AI to provide all solutions to the detriment of their learning. This echoes the observations of Fernandez and Cornell  (Fernandez and Cornell, 2024), who emphasized the need for careful integration of AI-driven code generation tools to avoid such overreliance. Help-seeking is crucial for students to grasp new concepts, acquire skills, and tackle challenges in their computing courses (Hou et al., 2024). However, when participants use full descriptions of programming tasks as prompts and accept complete Generative AI-generated answers, Generative AI may not effectively fulfill its role as a help-seeking tool that constructively aids struggling students, but an oracle that does the learners’ job for them. This concern was also raised by Jošt et al. (Jošt et al., 2024), highlighting the need for instructional strategies that emphasize breaking down problems and leveraging Generative AI for incremental learning. Our findings, alongside those of Prather et al.  (Prather et al., 2024), suggest that some students may struggle with new metacognitive difficulties when using generative AI tools, including being conceptually behind in course material but unaware of it due to a false sense of confidence. Further exploration is needed to understand the underlying reasons for this undesired behavior and to encourage a more constructive use of Generative AI that promotes deep understanding and problem-solving skills in computing education.
Our observations reveal two main types of behavior: iterative and linear. Iterative users refined their prompts to achieve correct answers, while linear users used full descriptions to find answers directly. This behavioral split reflects the findings of Vadaparty et al. (Vadaparty et al., 2024), who noted similar patterns in student interactions with Generative AI in a CS1 course. The iterative approach can improve learning by encouraging deeper engagement with problem-solving processes, whereas the linear approach may indicate a tendency to seek quick fixes.
Some prompts were related to coding concepts, indicating that novices struggle the most to understand programming concepts. This is consistent with the findings of Liu et al. (Liu et al., 2024), who found that students often used Generative AI tools to clarify the concepts of coding and the logic of the program. Our study further suggests that while Generative AI can assist with the coding concepts, additional instructional support is needed for the logic and debugging of the program.
Our results show that in 30% of prompts, users accepted the Generative AI response, while in 70%, AI responses were used to learn concepts. This dual role of Generative AI as both an answer provider and a personal tutor aligns with the observations by Prather et al. (Prather et al., 2024) regarding the mixed impact of Generative AI on student learning, where Generative AI facilitated both understanding and dependency. Effective Generative AI integration should balance assistance with promoting independent problem-solving skills.
We can categorize the nature of interactions with Generative AI based on the reliance of the prompt on the prior prompts. into two groups (1) stateless and (2) stateful. In the stateless interactions, individual prompts are independent of the prior prompts, hence prompts can be interpreted and answered independently. However, in stateful interactions, the prompt assumes that the Generative AI uses the history of the user’s interaction. For instance, P7 used stateful interaction, where he asked “So, what was wrong with my initial code?”.
We observed that seven participants P4, P7, P9, P10, P13, P14, and P15 employed the same approach on all tasks. We call their strategy for problem-solving, single-approach strategy. In contrast, P1, P2, P3, P5, P6, P8, P11, and P12 used different approaches for different tasks that we call mixed-approach.
As generative AI becomes more commonplace and more learners can access them for diverse sets of topics, a more granular investigation of these approaches becomes more important to guide pedagogy.
6.2. Self Efficacy and Generative AI
Our observations suggest that users who used full descriptions of questions may prefer smooth interactions without challenges. This behavior correlates with lower self-efficacy scores, similar to the findings of Xue et al. (Xue et al., 2024), who observed that students with lower self-efficacy tend to rely more heavily on Generative AI tools. Overall, most users’ self-efficacy levels increased after programming with Generative AI. This suggests the varied impacts of integrating Generative AI in educational contexts, shaping self-efficacy outcomes according to individual learning strategies and initial confidence levels. Enhancing self-efficacy through scaffolded Generative AI interactions could support more confident and independent problem-solving.
6.3. Implications for Computing Education
The current approach to prompt engineering focuses on productivity, aiming to help developers find final solutions quickly. However, in an educational context, the goal is not merely to reach a solution, but to ensure that students achieve the learning objectives. This approach is inherently different from current practices in prompt engineering, as proper a learning strategy may require expecting or even encouraging a significant amount of purposeful failures along the way, while productivity aims to minimize and eradicate those failures. Thus, it is necessary to investigate prompt engineering approaches that prioritize individual learning.
Our results suggest that, in the physical presence of a researcher and with knowledge of being recorded, in a considerable number of cases, the participants directly resorted to Generative AI to provide solutions to the tasks without making any attempts on their own. This undesirable way of help-seeking, if used as the default approach in solving homework problems, can negatively impact students’ learning. This suggests the need for students to improve their self-regulation skills, e.g. self-monitoring (Schunk and Zimmerman, 2012) to monitor and reflect the intensity and frequency of their use of Generative AI. Similarly, Generative AI educational tool builders should consider supporting such strategies that enhance students’ self-regulation and help-seeking behavior when they use Generative AI.
One can imagine that a possible remedy for this problem is to devise novel homework assignments that can be difficult for Generative AI to solve. However, as Generative AI tools become more sophisticated and powerful, the arms race between educators and Generative AI seems to be a losing proposition, especially in introductory courses such as CS1 that only include basic algorithmic thinking that there is an upper limit to the appropriate complexity of problems. To discourage students from using prompts to find complete solutions to homework problems, we should investigate novel types of problems that are compatible with the era of Generative AI. Recent work such as Promptly (Denny et al., 2024a) that uses the graphical representation for problem description instead of a textual description is a step in this direction.
An intriguing observation from our study was that most participants did not frequently execute their programs during development to verify the correctness or identify syntax errors. This lack of regular testing suggests a gap in their understanding of the importance of iterative debugging in the programming process. Consequently, students may miss out on early detection of mistakes, which can lead to more complex issues and frustration later in the development cycle. Addressing this behavior through new pedagogical tools such as (Ferdowsi et al., 2024) that visualizes the value of variables as students develop their solutions could improve students’ coding practices and overall comprehension of programming concepts.
Acknowledgments
This material is based upon work supported by the U.S. National Science Foundation under Grant No. 2225373. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.